OpenStudy (anonymous):

Consider line A whose equation is y = -10x + 8. If line B is perpendicular to line A and has a y-intercept at y = 3, what is the equation for line B?

OpenStudy (anonymous):

y=-10x+3 y=10x+3 y=1/10x+3 y=-1/10x+3

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@Mchilds15

OpenStudy (mchilds15):

Okay, this is super easy. But can you tell me what you think it is?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

brb

OpenStudy (mchilds15):

Actually, i would say it is c. If the lines are perpendicular, the slope for each line would be polar opposites. So line A has a slope of -10. The opposite of that would be 1/10 slope.

OpenStudy (bee_see):